OpenStudy (yanasidlinskiy):

Multiply or divide. Solve the equation. 5x + 10 x 2x x + 2 4x - 10

OpenStudy (anonymous):

@YanaSidlinskiy are the ones in the bottom the choices?

OpenStudy (yanasidlinskiy):

no its 5x +10/ x+2 multiply by 2x/4x-10

OpenStudy (tkhunny):

Have you studied Order of Operations? It may also be called "Precedence".

OpenStudy (yanasidlinskiy):

yeah...but i still can't figure it out...im like really stuck..

OpenStudy (tkhunny):

Then you should review that material. It will help A LOT to write clearly what you mean. 5x +10/ x+2 means \(5x + \dfrac{10}{x} + 2\) 2x/4x-10 means \(\dfrac{2x}{4x} - 10\) or maybe \(\dfrac{2x}{4}x - 10\) If you mean NONE of that, please use parentheses to clarify meaning. I suppose you may have meant: \(\dfrac{5x+10}{x+2}\cdot \dfrac{2x}{4x-10}\) Factor what there is to be factored. \(\dfrac{5(x+2)}{x+2}\cdot \dfrac{2x}{2(2x-5)}\) Rearrange a bit. \(5\cdot\dfrac{x+2}{x+2}\cdot\dfrac{2}{2}\cdot\dfrac{x}{2x-5}\) Simplify where you can. \(5\cdot\dfrac{x}{2x-5} = \dfrac{5x}{2x-5}\) Seriously, try to use proper notation. It WILL save you!

OpenStudy (yanasidlinskiy):

Is the answer gonna be like \[\frac{ 10x+10 }{ 4x + 20}\]

OpenStudy (tkhunny):

Maybe something like that, but there is more to do.